A railway track runs parallel to a road until a turn brings
the road to railway crossing. A cyclist rides along the road
everyday at a constant speed 20 kmh. He normally meets
a train that travels in same direction at the crossing. One
day he was late by 25 min and met the train 10 km before
the railway crossing. Find the speed of the train.​

Grade:12

1 Answers

Arun
4 years ago
the cyclist was 10/20 = half an hour away from the crossing the moment he met the train on the "late day". on a usual day he'd reach the crossing in just 30-25=5 minutes. it means the train had to move 10 miles in 5 minutes to "meet" the cyclist on a usual day at the crossing. 

train's speed 
= 10 km / 5minutes 
= 10 / (5/60) 
= 120 kmph